The members of the band were selling t-shirts and school banners for a fundraiser. Justine sold 13 tshirts and 10 banners for a total of $145.00. Carlos sold 15 tshirts and 6 banners for a total of $141.00. How much did the band charge for tshirts? How much for banners?
Answer by mananth(16949)

let T shirts cost $x
Let banners cost y
...
13x+10y=145
15x+6y= 141
.....
13x+10y =145 .............1
15x+6y= 141 .............2

multiply (1)by 3
Multiply (2) by -5
39x+30y =435
-75x-30 y=-705

Add the two equations

-36x=-270
/-36

x= $7.5

plug value of x in (1)

13x+10y =145
97.5+10y=145
10y=145 -97.5
10y=47.5
y= $4.75